OpenStudy (anonymous):

Write the equation of the line, in standard form, that passes through the origin and is parallel to x + y = 6

OpenStudy (anonymous):

@sleepyjess @hartnn

OpenStudy (sleepyjess):

Ok, what is the slope of the given line?

OpenStudy (anonymous):

Slope is -1 and y-intercept is 6

OpenStudy (anonymous):

Good.. :)

OpenStudy (sleepyjess):

Yes, so we know that the slope of the unknown line is -1, and it passes through the origin, so the y-intercept is 0.

OpenStudy (anonymous):

So how would this one be set up?

OpenStudy (sleepyjess):

Well, we would use slope intercept form first again

OpenStudy (anonymous):

y=-1x+0 ..?

OpenStudy (sleepyjess):

Yes, now move the -x to the other side

OpenStudy (anonymous):

-x+y=0 but multiply by -1 so x-y=0. But where does the 6 go?

OpenStudy (anonymous):

Be careful..!! :P

OpenStudy (sleepyjess):

We need to do the opposite

OpenStudy (anonymous):

Now what? Lol

OpenStudy (anonymous):

you will add \(x\) both the sides and not subtract..

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@eta What do I need to do?

OpenStudy (anonymous):

\(y = -x\)

OpenStudy (anonymous):

Add \(x\) dear..!!

OpenStudy (anonymous):

x+y=0

OpenStudy (anonymous):

\[y + x = -x + x \implies x+y = 0\]

OpenStudy (anonymous):

Good, up to the mark..

OpenStudy (anonymous):

But what do I need to do with the 6?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Write the equation of the line, in standard form, that passes through the origin and is parallel to x + y = 6

OpenStudy (anonymous):

See, there are two equations, or you can say you have two lines..

OpenStudy (anonymous):

There can't be two equations

OpenStudy (anonymous):

One line you are totally given, ie x+y =6 Now, you are finding the equation of another line which is parallel to this line..

OpenStudy (anonymous):